Mr Zoo is planning to set up a mini Zoo in Howick.

The following relates to different pricing plans of visiting the zoo and costs of running the zoo.

Pricing Plan 1 - $30 per person (a special pack will be included)
Pricing Plan 2 - $25 per visitor (the special pack not to be included)

Variable Costs per visitor

Special Pack - $8
Zoo Consumables - $7
Zoo Expenses - $2

Total Fixed Costs - $200,000


REQUIRED: Study the information above then answer the following questions.

Using the CVP formula, calculate the number of units (and its dollar amount) for achieving
Breakeven
A profit of $30,000 for both plans.
Calculate the Contribution Margin ratio for both plans.
Describe the meaning of Contribution Margin ratio.
Explain the relationship between Contribution Margin ratio and Number of Breakeven units.
Explain why it is important to calculate Margin of Safety.
1
Expert's answer
2018-08-05T07:15:08-0400
(a) a. Using he CVP formula, we can compute the Breakeven Point in units:
Plan 1:
x = 200 000 / (30 - (8 + 7 + 2) = 15 385 units
Break-Even Point in Sales Revenue:
200 000 / ((30 - (8 + 7 + 2)) / 30) = $352 941,2
Plan 2:
x=200 000 / (25 - (7 + 2) = 12 500 units
Break-Even Point in Sales Revenue
200 000 / ((25 - (7 + 2)) / 25) = $312 500
b. A profit of $30,000 for both plans.
Plan 1:
x = (200 000 + 30 000) / (30 - (8 + 7 + 2) = 17 692 units
Sales Revenue:
(200 000 + 30 000) / ((30 - (8 + 7 + 2)) / 30) = $405 882,4
Plan 2:
x = (200 000 + 30 000) / (25 - (7 + 2) = 14 375 units
Sales Revenue:
(200 000 + 30 000) / ((25 - (7 + 2)) / 25) = $4638 888, 9

(b) Calculate the Contribution Margin ratio for both plans.
Plan 1: (30 - (8 + 7 + 2)) / 30) = 0,43 = 43%
Plan 2: (25 - (7 + 2)) / 25) = 0,64 = 64%

(с) According to Plan 2, the Contribution Margin ratio is higher (64% > 32%), which means greater efficiency.

(d) in both formulas, the difference between price and variable costs is used in calculations.

(e) the calculation of the safety margin is important for determining the risks of activity.